Shift Summary Day 1 Aug., 1999 ('fete national' Switzerland)

Nickolas Solomey (solomey@hep.uchicago.edu)
Sun, 1 Aug 1999 15:30:43 -0500

Shift Summary for day shift, 08/1/99

On Shift: Nick, Doug J.

Hours of beam available at all intensities: 8.5
Hours of beam available at nominal intensity: 0
(however they were close ~3.8e12 instead of 4.2e12 would be 8.5 hrs
or 7.69 effective hours at nominal intensity)
Hours of E832 data (special runs not included): 8.25
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: 12 min changing tapes, several spills due to
kfrend.
========================================================================
Tapes brought to FCC this shift:QKa20y (for a=C,S,K and y=1,2,3), QKB069.
========================================================================
Accesses during shift: none
========================================================================
Overview of shift: Took data!
========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14077 685 7:00 Take over shift, DA is running with good data.
14077 711 7:32 Automatic End-of-Run.
14078 1 7:45 Start new run. Switch about as fast as possible!
7:47 Respond to request for split tape mount.
KQN017 mounted.
157 11:15 Val message - laser too wide - disappeared.
165 11:25 Triggers stopped - KFRIENDed it - they started
KFRIENDed a few more times at "Elliot's request"
11:55 SDAQ error, paged Ashkan who said it was not
serious but he would be in shortly to take care
of it.
12:25 SDAQ reset remotely by Ashkan.
12:30 Run manager calls, was surpised everything was
good, next time Doug suggest we make something up.
221 12:41 KTeV ana errors, so we ran KFREND.
13:45 Have mounted KQNL16 on KTeV1 as per request.
13:55 Dave Northacker filled the drift chamber gas
system with alchole.
14:05 Got a couple indications of ktsad.exe not running
paged Ashkan. problems do not recur
314 14:45 KFRIEND 3x
323 Val message - laser two wide - disappeared.
343 15:22 Val message again - laser two wide - disapperaed.
348 15:30 Turned over run to next shift.

=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=
Dr. Nickolas Solomey, Research Scientist, Enrico Fermi Institute
The University of Chicago, 5640 S. Ellis Ave, Chicago, Ill. 60637 USA
Phone: (773) 702-3278, fax: x-6317, http://hep.uchicago.edu/~solomey